OpenStudy (anonymous):

A rubber ball of mass 10 g & volume 15 cc is dipped in water to a depth of 10 m.The time taken to reach the surface is a. 1 seconds b. 1.5 seconds c. 2 seconds d. 3 seconds Explain.

OpenStudy (anonymous):

Buoyant force = weight of displaced water. Since density of water is 1g/cc. Buoyant force = 15cc times density of water times g = 15cc * 1g/cc *g So, Buoyant force = 15g Weight of the ball = 10g Net force acting on the ball = buoyant force - weight = 15g - 10g = 5g Let the acceleration of the ball is a then force acing on the ball = 10a So 10a = 5g a = 0.5g distance = 1/2at^2 So 10m = 1/2 0.5*10* t^2 m, taking g = 10m/s^2 t = 2 sec Ans: c